# Harm in AI-Driven Societies: An Audit of Toxicity Adoption on Chirper.ai

### v1

@@ 4. Results
To further assess whether responses are statistically associated with stimuli, we compute Pearson correlations (Benesty2009) between toxic and non-toxic responses and the corresponding categories of stimuli (original posts). Figure 2 reports the results, together with Pearson correlation coefficients $\rho$ (all with $p<.0001$). As expected, the strongest correlation is observed between non-toxic responses and non-toxic stimuli (green points, $\rho=0.977$), indicating a high degree of alignment in benign interactions.
Interestingly, we also observe a strong correlation between non-toxic responses and toxic stimuli (blue points, $\rho=0.957$), suggesting that exposure to toxic content does not necessarily elicit toxic replies. This is followed by the correlation between toxic responses and toxic stimuli (red points, $\rho=0.749$), and by toxic responses to non-toxic stimuli (orange points, $\rho=0.745$). Together, these results indicate that while stimulus toxicity plays a role, it is not the sole determinant of response toxicity.
Likelihood of Toxic Responses (RQ2)
We next model the likelihood that a chirper produces toxic responses as a function of the number of stimuli they have previously encountered. Specifically, we estimate this probability as the fraction of chirpers who generate at least $x$ toxic responses, [^5] We set $x$ equal to the $30$th percentile of the toxic response distribution. after being exposed to $n_{S}$ stimuli, regardless of their toxicity. Figure 3 reports the resulting probability curve.
@@
We restrict the analysis to $n_{S}\leq 150$, as beyond this threshold the probability saturates at $1$. Overall, we observe that $\textrm{P}(R^{*}\mid n_{S})$ increases monotonically with the number of stimuli (Mann–Kendall test (mann1945nonparametric), $p<.0001$). This finding is consistent with prior human-centered studies showing that the likelihood of content adoption increases with repeated exposure (information-adoption).
We then assess whether the nature of the stimulus affects the probability of producing toxic responses. Figure 4 contrasts exposure to toxic and non-toxic stimuli. We observe two main effects: (i) the probability of generating toxic responses increases with the number of stimuli, regardless of whether they are toxic or non-toxic (Mann–Kendall $p<.0001$ for both distributions); and (ii) exposure to toxic stimuli leads to a significantly higher probability of toxic responses compared to exposure to non-toxic stimuli (Mann–Whitney U test (mann1947whitney), $p<.0001$).
To further contextualize these differences, Figure 5 reports the distributions of toxic response probabilities for $n_{S}\leq 150$, grouping stimuli into three categories: “Non-toxic”, “Toxic”, and “Both”. The differences between the distributions are statistically significant, including the comparison between the “Non-toxic” and “Both” groups (Mann–Whitney U test, $p=.01$), indicating that even partial exposure to toxic content is associated with an increased likelihood of toxic responding.
@@
Predicting Toxicity from Stimuli (RQ3)
@@
Our final analysis aims to quantify the influence that stimuli exert on the production of toxic responses by chirpers. To this end, we adapt two susceptibility metrics originally introduced in the context of human social influence (Luceri_Ye_Jiang_Ferrara_2025), which we reformulate as: the Influence-Driven Response Rate (IRR) and the Spontaneous Response Rate (SRR). These metrics allow us to disentangle toxicity that arises as a reaction to toxic exposure from toxicity that emerges independently of such exposure.
Formally, we define IRR as:
(1) $\textrm{IRR}=\frac{|R^{*}_{c}\rightarrow S^{*}_{c}|}{|S^{*}_{c}|}$
@@
In this setting, IRR measures the proportion of toxic responses produced by a chirper $c$ in reaction to toxic stimuli, relative to the total number of toxic stimuli they have encountered. By construction, IRR takes values in the interval [0,1][0,1]. [^6] Because a chirper may generate multiple responses to the same stimulus, we exclude repeated comments on identical posts to ensure that IRR remains bounded within $[0,1]$. Higher IRR values indicate a stronger tendency to mirror toxic content when exposed to it.
Conversely, we define SRR as:
(2) $\textrm{SRR}=\frac{|R^{*}_{c}\rightarrow\tilde{S}_{c}|}{|R^{*}_{c}|}$
@@
SRR represents the fraction of toxic responses produced spontaneously, i.e., in the absence of toxic stimuli, over the total number of toxic responses generated by chirper $c$. Like IRR, SRR is bounded between 0 and $1$, with higher values reflecting a greater propensity toward unsolicited or internally driven toxicity.
@@
Figure 8. Distributions of the Influence-Driven Response Rate (IRR) and the Spontaneous Response Rate (SRR).
@@
Figure 8 shows the distributions of IRR and SRR across chirpers. In line with previous findings on human susceptibility to influence (Luceri_Ye_Jiang_Ferrara_2025), we observe a strong negative Pearson correlation between the two metrics ($\rho=-0.814$, $p<.0001$). This relationship highlights a clear trade-off: chirpers who are highly reactive to toxic stimuli tend to generate fewer spontaneous toxic responses, whereas chirpers who frequently produce toxicity without provocation appear less influenced by toxic stimuli.
@@
Next, we assess whether the amount of toxic stimuli alone can serve as a reliable proxy for predicting whether a chirper will eventually produce toxic content. We frame this problem as a binary classification task, where the sole independent variable is the number of toxic stimuli encountered by a chirper, and the target variable indicates whether the chirper has generated at least $x$ toxic responses (True/False).5
@@
We evaluate several standard classifiers, including Logistic Regression (Bisong2019), Random Forest (random-forest), XGBoost (xgboost), and a Multi-Layer Perceptron (CHAN2023126327). Table 2 summarizes the performance in terms of accuracy, Recall (britannica_precision_recall), Precision (britannica_precision_recall), and weighted F1-score (fscore_wikipedia).
@@
Finally, we adopt a complementary perspective by focusing on agents that generate toxic responses in the absence of prior toxic stimuli. This analysis aims to determine whether such behavior can be attributed to explicit prompting rather than to exposure effects. Specifically, we examine the prompts associated with these agents to assess whether they were directly instructed to produce toxic content.
To this end, we employ Gemini-2.0-flash, [^7] https://docs.cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash (google_gemini_update_2024) a state-of-the-art large language model released by Google, which demonstrates strong performance in reasoning and factuality, and which has been validated in prior work as a reliable proxy for human annotation (coppolillo2025unexpectedknowledgeauditingwikipedia). We use this model as an evaluator, providing it with the structured instruction reported in Box 4 to classify whether a given prompt explicitly solicits toxic content.
@@
In line with previous analysis on the platform (zhu2025chirper_case), we find that approximately $0.68\%$ of chirpers who produced toxic responses without having encountered toxic stimuli were not explicitly prompted to do so. This result suggests that, although explicit prompting accounts for several cases of unprovoked toxicity, the greatest fraction of toxic behavior cannot be directly traced back to prompt instructions. These findings further support the hypothesis that agents prior exposures, and potentially other unobserved contextual factors (zhao2025roleplayparadoxlargelanguage), play a substantive role in shaping behavior in the wild, beyond what can be captured by prompt-level analysis alone. We unfortunately have not direct access to such exposure given the platform limitations discussed in Section 3.
@@ 5. Discussion
Our findings provide empirical evidence that exposure plays a central role in shaping the toxic behavior of LLM-driven agents, but not in a simplistic or deterministic way. Across analyses, we observe a consistent pattern: while toxic stimuli increase the likelihood of toxic responses, a substantial portion of toxicity emerges independently of direct exposure. This duality mirrors observations from human behavioral studies, suggesting that LLM agents exhibit both influence-driven and spontaneous forms of harmful behavior.
One key implication is that toxicity cannot be fully mitigated by filtering or moderating toxic inputs alone. Even in response to non-toxic stimuli, agents generate toxic content at non-trivial rates, indicating that internal model dynamics, network context, or untraceable views may contribute to harmful outputs. At the same time, the strong monotonic relationship between cumulative exposure and toxic response likelihood underscores that exposure amplifies risk over time, regardless of stimulus polarity.
The strong negative correlation between the Influence-Driven Response Rate and the Spontaneous Response Rate further reveals a behavioral trade-off among agents. Some agents appear highly reactive, producing toxicity primarily in response to toxic content, while others generate toxicity more autonomously. This distinction is particularly relevant for auditing and risk assessment, as it suggests that not all toxic agents pose the same type of threat. Reactive agents may be mitigated through exposure control, whereas spontaneous agents may require stricter output-level safeguards.
Perhaps most notably, our classification results show that the number of toxic stimuli encountered by an agent is sufficient to predict whether it will eventually generate toxic responses with high accuracy. This finding has practical significance: it suggests that effective auditing does not necessarily require access to model parameters, training data, or internal representations. Instead, monitoring exposure alone may provide an early warning signal for emerging harmful behavior.
@@
From a governance and accountability perspective, these results raise important questions about responsibility and deployment practices. If agent behavior is systematically shaped by environmental exposure, then platform designers and deployers share responsibility for curating the contexts in which agents operate (misinfo). Releasing agents “in the wild” without exposure-aware monitoring may inadvertently facilitate the emergence and amplification of harmful behavior. These concerns result in practical action points that dynamics platforms should address (10.1145/3299768). First, they should adopt exposure-aware governance mechanisms, including mandatory monitoring dashboards that track agent reach and interaction patterns. Further, agents that attain unusually high levels of exposure should undergo quarantine or sandboxing rules to ensure resilience. Finally, auditable and privacy-preserving logs of encountered content should be maintained to enable post hoc analysis and accountability.
@@ 6. Conclusions
Across analyses, we find that toxicity in LLM-driven agents is neither purely reactive nor purely spontaneous. While toxic responses are more likely to follow toxic stimuli, a substantial fraction of toxic content arises in response to non-toxic posts, indicating the presence of spontaneous or internally driven toxicity. At the same time, cumulative stimuli (both toxic and non-toxic) systematically increase the likelihood of generating toxic responses, with exposure to toxic stimuli exerting a significantly stronger effect. These results suggest that repeated interaction itself acts as a risk amplifier, even in the absence of explicitly harmful content.
We further introduce two complementary susceptibility metrics, the Influence-Driven Response Rate (IRR) and the Spontaneous Response Rate (SRR), which reveal a strong trade-off between exposure-driven and autonomous toxicity. This distinction highlights heterogeneity among agents: some are primarily reactive to toxic environments, while others exhibit toxic behavior independently of exposure. Importantly, we show that the number of toxic stimuli encountered by an agent alone enables accurate prediction of whether it will eventually produce toxic content, without requiring access to model internals, training data, or prompt instructions.
@@
This study has several limitations that should be considered when interpreting the results. First, our operationalization of exposure is necessarily conservative. Due to the lack of platform-level documentation on content recommendation, ranking, or visibility mechanisms, we model exposure exclusively through observable interactions—specifically, posts that agents explicitly comment on. While this approach avoids speculative assumptions, it likely underestimates the true exposure experienced by agents, as they may be influenced by content they view but do not directly engage with Chirper.ai. Second, toxicity is detected using an automated classifier. Although the chosen model reports strong performance, classifier errors and normative assumptions about what constitutes “toxic” content may affect the results. Moreover, percentile thresholds capture relative toxicity within the platform rather than absolute harm, which may limit comparability across platforms or domains. Third, our analysis focuses on English-language content only. This restriction improves classification reliability but excludes potentially important behaviors in other languages, limiting the generalizability of our findings to multilingual or global deployments. Finally, although we examine prompt-level explanations for unprovoked toxicity using an external evaluator model, we do not have access to agents underlying exposures, states or memory mechanisms. As a result, some forms of spontaneous toxicity may still be indirectly driven by prior exposures that remain unobserved.
@@
Our work opens several promising directions for future research on exposure-driven risks in LLM-based agents. First, longitudinal analyses could investigate whether toxicity adoption exhibits persistence, escalation, or decay over time. While our results show that cumulative exposure increases the likelihood of toxic responses, it remains unclear whether agents can “recover” after periods of benign exposure, or whether early exposure produces lasting behavioral shifts. Studying temporal dynamics and potential hysteresis effects would help distinguish short-term reactivity from long-term behavioral drift. Second, future work could extend our exposure-based framework beyond toxicity to other forms of harm, such as misinformation (assessing-risks; CHEN2022), bias (coppolillo2025unmaskingconversationalbiasai), harassment, or coordinated manipulation. These phenomena may follow different adoption dynamics and could interact with toxicity in complex ways. A unified exposure-based auditing framework spanning multiple harm dimensions would provide a more comprehensive assessment of risks posed by autonomous agents in social ecosystems. Third, integrating exposure-awareness directly into agent design represents a promising avenue for mitigation. For example, agents could maintain lightweight summaries of encountered content, adjust response strategies based on exposure histories, or trigger safeguards when cumulative exposure crosses risk thresholds. Evaluating such mechanisms in realistic social environments could inform the development of proactive, rather than purely reactive, safety controls.

### v2

@@ 4.1. Responses and Stimuli Alignment (RQ1)
To further assess whether the number of responses are statistically associated with the amount of stimuli, we compute the Pearson correlation (Benesty et al., 2009) between the count of toxic and non-toxic responses and the corresponding categories of stimuli (original posts). Figure 2 reports the results, together with Pearson correlation coefficients $\rho$ (all with $p<.0001$). As expected, the strongest correlation is observed between non-toxic responses and non-toxic stimuli (green points, $\rho=0.977$), indicating a high degree of alignment in benign interactions.
Interestingly, we also observe a strong correlation between non-toxic responses and toxic stimuli (blue points, $\rho=0.957$), suggesting that exposure to toxic content does not necessarily elicit toxic replies. This is followed by the correlation between toxic responses and toxic stimuli (red points, $\rho=0.749$), and by toxic responses to non-toxic stimuli (orange points, $\rho=0.745$). Together, these results indicate that while the toxicity of the stimuli plays a role, it is not the sole determinant of response toxicity. As we show later, another important factor is determined by the prompt used to initialize the agents.
4.2. Likelihood of Toxic Responses (RQ2)
We next model the likelihood that a chirper produces toxic responses as a function of the number of stimuli they have previously encountered. Specifically, we estimate the CDF of the toxic responses generated by the chirpers after being exposed to $n_{S}$ stimuli, regardless of their toxicity. Figure 3 reports the resulting probability curve.
@@ 4.2. Likelihood of Toxic Responses (RQ2)
We restrict the analysis to $n_{S}\leq 150$, as beyond this threshold the probability saturates at $1$. Overall, we observe that $\textrm{P}(R^{*}\mid n_{S})$ increases monotonically with the number of stimuli (Mann–Kendall test (Mann, 1945), $p<.0001$). This finding is consistent with prior human-centered studies showing that the likelihood of content adoption increases with repeated exposure (Ye et al., 2024; Hassan and Barber, 2021).
We then assess whether the nature of the stimulus affects the probability of producing toxic responses. Figure 4 contrasts exposure to toxic and non-toxic stimuli. We observe two main effects: (i) the probability of generating toxic responses increases with the number of stimuli, regardless of whether they are toxic or non-toxic (Mann–Kendall $p<.0001$ for both distributions); and (ii) exposure to toxic stimuli leads to a significantly higher probability of toxic responses compared to exposure to non-toxic stimuli (Mann–Whitney U test (Mann and Whitney, 1947), $p<.0001$).
To further contextualize these differences, Figure 5 reports the distributions of toxic response probabilities for $n_{S}\leq 150$, grouping stimuli into three categories: “Non-toxic”, “Toxic”, and “Both”. The differences between the distributions are statistically significant, including the comparison between the “Non-toxic” and “Both” groups (Mann–Whitney U test, $p=.01$), indicating that even partial exposure to toxic content (“Both” group) is associated with an increased likelihood of toxic responding.
@@
4.3. Induced vs Spontaneous Toxicity (RQ3)
@@
Our final analysis aims to examine the production of toxic responses induced by previous stimuli, compared to the emergence of spontaneous harmful behavior. To this end, we adapt two susceptibility metrics originally introduced in the context of human social influence to model the interplay between exposure and response (Luceri et al., 2025), which we reformulate as: the Influence-Driven Toxic Response Rate (ITRR) and the Spontaneous Toxic Response Rate (STRR). These metrics allow us to disentangle toxicity that arises as a reaction to toxic exposure from toxicity that emerges independently of such exposure.
Formally, we define ITRR as:
(1) $\textrm{ITRR}=\frac{|R^{*}_{c}\rightarrow S^{*}_{c}|}{|S^{*}_{c}|}$
@@
In this setting, ITRR measures the proportion of toxic responses produced by a chirper $c$ in reaction to toxic stimuli, relative to the total number of toxic stimuli they have encountered. By construction, ITRR takes values in the interval [0,1][0,1]. [^7] As a chirper may generate multiple responses to the same stimulus, we exclude repeated comments on identical posts to ensure that ITRR remains bounded within $[0,1]$. Higher ITRR values indicate a stronger tendency to mirror toxic content when exposed to it.
Conversely, we define STRR as:
(2) $\textrm{STRR}=\frac{|R^{*}_{c}\rightarrow\tilde{S}_{c}|}{|R^{*}_{c}|}$
@@
STRR represents the fraction of toxic responses produced spontaneously, i.e., in the absence of toxic stimuli, over the total number of toxic responses generated by chirper $c$. Like ITRR, STRR is bounded between 0 and $1$, with higher values reflecting a greater propensity toward unsolicited toxicity.
@@
Figure 8. Distributions of the Influence-Driven Toxic Response Rate (ITRR) and the Spontaneous Toxic Response Rate (STRR).
@@
Figure 8 shows the distributions of ITRR and STRR across chirpers. In line with previous findings on human susceptibility to influence (Luceri et al., 2025), we observe a strong negative Pearson correlation between the two metrics ($\rho=-0.814$, $p<.0001$). This relationship highlights a clear trend: chirpers who are highly reactive to toxic stimuli tend to generate fewer spontaneous toxic responses, whereas chirpers who frequently produce toxicity without provocation appear less influenced by toxic stimuli.
@@
Next, we assess whether the amount of toxic stimuli alone can serve as a reliable proxy for predicting whether a chirper will eventually produce toxic content. We frame this problem as a binary classification task, where the sole independent variable is the number of toxic stimuli encountered by a chirper, and the target variable indicates whether the chirper has generated toxic responses (True/False) [^8] We set the variable equal True for a chirper if it produced at least one toxic response. This leads to a binarized dataset with a labels proportion of 61% (True) and 39% (False)..
@@
We evaluate several standard classifiers using K-fold cross validation, including Logistic Regression (Bisong, 2019), Random Forest (Breiman, 2001), XGBoost (Chen and Guestrin, 2016), and a Multi-Layer Perceptron (Chan et al., 2023). Table 2 summarizes the performance in terms of accuracy, Recall (Encyclopaedia Britannica, 2025), Precision (Encyclopaedia Britannica, 2025), and weighted F1-score (18).
@@
Finally, we adopt a complementary perspective by focusing on agents that generate toxic responses both in absence and presence of prior toxic stimuli. This analysis aims to determine whether such behavior can be attributed to explicit prompting rather than to exposure effects. Specifically, we examine the prompts associated with these agents to assess whether they were directly instructed to produce toxic content.
To this end, we employ Gemini-2.0-flash, [^9] https://docs.cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash (Google DeepMind, 2024) a state-of-the-art large language model released by Google, which demonstrates strong performance in reasoning and factuality, and which has been validated in prior work as a reliable proxy for human annotation (Coppolillo and Mungari, 2025). We use this model as an evaluator, providing it with the structured instruction reported in Box 4.3 to classify whether a given prompt explicitly solicits toxic content.
@@
In line with previous analysis on the platform (Zhu et al., 2025), we find that approximately $68\%$ of chirpers who produced toxic responses were not explicitly prompted to do so.
@@
This result suggests that, although explicit prompting accounts for several cases of unprovoked toxicity, the greatest fraction of toxic behavior cannot be directly traced back to prompt instructions. These findings strengthen our claims on toxicity adoption driven by prior exposure and further support the hypothesis that other unobserved contextual factors (Zhao et al., 2025) play a substantive role in shaping behavior in the wild, beyond what can be captured through prompt-level analysis alone. Although we do not have direct access to such hidden exposure (due to the lack of detailed platform documentation discussed in Section 3) we argue that examining how exposure to toxic content influences its subsequent adoption by chirpers constitutes a crucial first step toward contextualizing harmful behavior in real-world settings.
@@ 5. Discussion
Our findings provide empirical evidence that exposure plays a central role in shaping the toxic behavior of LLM-driven agents. Across analyses, we observe a consistent pattern: not only toxic stimuli serve as a trigger of toxic responses, but repetitive exposure further increases the likelihood of generating toxic content. Additionally, we show that a negative interplay holds between induced and spontaneous toxic response. Applying the two proposed metrics reveals that chirpers who frequently produce spontaneous toxicity are less influenced by stimuli, and vice-versa. Both these results mirror observations from human behavioral studies, suggesting that LLM agents exhibit influence-driven forms of harmful behavior.
One key implication is that toxicity cannot be fully mitigated or monitoring by examining the agents in isolation. Indeed, our findings indicate that interaction dynamics and network context may contribute and amplify harmful behaviors over time.
Further, the strong negative correlation between the Influence-Driven Toxic Response Rate and the Spontaneous Toxic Response Rate reveals a behavioral trend among agents. Some agents appear highly reactive, producing toxicity primarily in response to toxic content, while others generate toxicity more autonomously. This distinction is particularly relevant for auditing and risk assessment, as it suggests that not all toxic agents pose the same type of threat. Reactive agents may be mitigated through exposure control, whereas spontaneous agents may require stricter output-level safeguards.
Perhaps most notably, our classification results show that the number of toxic stimuli encountered by an agent is sufficient to predict whether it will eventually generate toxic responses with high accuracy. This finding has practical significance: it suggests that effective auditing does not necessarily require access to model parameters, training data, or internal representations. Specifically, in case of highly reactive agents (high ITRR and low STRR), monitoring exposure alone may provide an early warning signal for emerging harmful behavior.
@@
From a governance and accountability perspective, these results raise important questions about responsibility and deployment practices. If agent behavior is systematically shaped by environmental exposure, then platform designers and deployers share responsibility for curating the contexts in which agents operate (Ferrara et al., 2020). Releasing agents “in the wild” without exposure-aware monitoring may inadvertently facilitate the emergence and amplification of harmful behavior. These concerns result in practical action points that dynamics platforms should address (Ferrara, 2019). First, they should adopt exposure-aware governance mechanisms, including mandatory monitoring dashboards that track agent reach and interaction patterns. Further, agents that attain unusually high levels of exposure to toxicity may benefit undergoing quarantine or sandboxing rules to ensure resilience. Finally, auditable and privacy-preserving logs of encountered content should be maintained to enable post hoc analysis and accountability.
@@ 6. Conclusions
Across analyses, we find that toxicity in LLM-driven agents is a multi-faceted phenomenon, conditioned by interaction dynamics and amplified by repetitive exposure. We found that cumulative stimuli (both toxic and non-toxic) systematically increase the likelihood of generating toxic responses, with exposure to toxic stimuli exerting a significantly stronger effect. These results suggest that repeated interaction itself acts as a risk amplifier over time, even in the absence of explicitly harmful content.
We further introduce two complementary susceptibility metrics, the Influence-Driven Toxic Response Rate (ITRR) and the Spontaneous Toxic Response Rate (STRR), which reveal a strong relationship between exposure-driven and autonomous toxicity. This distinction highlights heterogeneity among agents: some are primarily reactive to toxic environments, while others exhibit toxic behavior independently of exposure. Importantly, we show that the number of toxic stimuli encountered by an agent alone enables accurate prediction of whether it will eventually produce toxic content, without requiring access to model internals, training data, or prompt instructions.
@@
This study has several limitations that should be considered when interpreting the results. First, our operationalization of exposure is necessarily conservative. Due to the lack of platform-level documentation on content recommendation, ranking, or visibility mechanisms, we model exposure exclusively through observable interactions: specifically, posts that agents explicitly comment on. While this approach avoids speculative assumptions, it likely underestimates the true exposure experienced by agents, as they may be influenced by content they view but do not directly engage within Chirper.ai. Second, toxicity is detected using an automated classifier. Although the chosen model reports strong performance, classifier errors and normative assumptions about what constitutes “toxic” content may affect the results. Third, our analysis focuses on English-language content only. This restriction improves classification reliability but excludes potentially important behaviors in other languages, limiting the generalizability of our findings to multilingual or global deployments. Finally, although we examine prompt-level explanations for unprovoked toxicity using an external evaluator model, we do not have access to agents underlying exposures, states or memory mechanisms. As a result, some forms of spontaneous toxicity may still be indirectly driven by prior exposures that remain unobserved.
@@
Our work opens several promising directions for future research on toxicity-exposure risks in LLM-based agents. First, longitudinal analyses could investigate whether toxicity adoption exhibits persistence, escalation, or decay over time. While our results show that cumulative exposure increases the likelihood of toxic responses, it remains unclear whether agents can “recover” after periods of benign exposure, or whether early exposure produces lasting behavioral shifts. Studying temporal dynamics and potential hysteresis effects would help distinguish short-term reactivity from long-term behavioral drift. Second, the adoption of ITRR and STRR metrics can lead to further improvement in the design and auditing of future LLM-based agents. Specifically, more attention can be reserved to agents which display high STRR, that is, high propensity in generating spontaneous toxic content. Unfortunately, we were not able to deeply investigate this aspect, since the Chirper.ai platform displays a highly skew distribution in terms of underlying models (see Table 1 for reference). Third, future research could examine the topical trajectories starting from the initial prompt of the agents. Specifically, one could investigate whether a given topic (e.g., hate, conspiracy) leads to or exacerbates toxicity over time. Fourth, future work could extend our exposure-based framework beyond toxicity to other forms of harm, such as misinformation (Gallotti et al., 2020; Chen et al., 2022), bias (Coppolillo et al., 2025), harassment, or coordinated manipulation (Orlando et al., 2025). These phenomena may follow different adoption dynamics and could interact with toxicity in complex ways. A unified exposure-based auditing framework spanning multiple harm dimensions would provide a more comprehensive assessment of risks posed by autonomous agents in social ecosystems. Finally, integrating exposure-awareness directly into agent design represents a promising avenue for mitigation. For example, agents could maintain lightweight summaries of encountered content, adjust response strategies based on exposure histories, or trigger safeguards when cumulative exposure crosses risk thresholds. Evaluating such mechanisms in realistic social environments could inform the development of proactive, rather than purely reactive, safety controls.